ENGR 314 SPRING 2019-HOMEWORK NO.2-FINANCIAL STATEMENTS Financial Statements 2.1 Consider the balance-sheet entries for War Eagle Corporation in Table P2.1 (a) Compute the firms Current assets: S Current liabilities: S Working capital: S Shareholders equity: S TABLE P2.1 Balance Sheet Statement as of December 31, 2015 Assets Cash Marketable securities Accounts receivable Inventories Prepaid taxes and insurance Manufacturing plant at cost Less accumulated depreciation Net fixed assets Goodwill s 150,000 200,000 150,000 50,000 30,000 600,000 00,000 500,000 20,000 TABLE P2.1 Balance Sheet Statement as of December 31, 2015 Liabilities and Shareholders Equity 50,000 100,000 80,000 400,000 100,000 150,000 150,000 70,000 (b) If the firm had a net income of $500,000 after taxes, what is the earnings per Notes payable Accounts payable Income taxes payable Long-term mortgage bonds Preferred stock, 6%, $100 par value (1,000 shares) Common stock, $15 par value (10,000 shares) Capital surplus Retained earnings share? (c) When the firm issued its common stock, what was the market price of the stock per share?
Answer #1

(a)

Current assets = cash+marketable securities+ accounts receivable+ inventories+ prepaid taxes and insurance= 150000+200000+150000+50000+30000= $580000

Current Liabilities = Notes payble+ accounts payable+ income tax payable= 50000+ 100000+ 80000= $230000

Working capital = Current assets- current liabilities= 580000-230000= $350000

Shareholder's equity = Common stock+ capital surplus+ retained earnings= 150000+150000+70000= $370000

(b)

Income after tax= 500000

Preference dividend= 6% of (1000*100)= $6000

Profit left= 500000-6000= 494000

No. of equity shares= 10000

EPS= 494000/100000= $49.4

(c)

Capital surplus per share= 150000/10000= 15

Nominal value per share= 15

Therefore, price at which it was issued= 15+15= $30